A cone is mounted on the centre of the flat surface of a hemi sphere of radius 14 cm. Radius of the cone is 7 cm. If the height of cone is 7 cm then Find surface area of it.
Answer by mananth(16949)

surface area of sphere = 4pir^2
surface area of hemisphere = 1/2 pir^2
r=14
surface area = 1/2*pi*(14^2)
=307.87 sq.cm
..
surface area of cone = pir^2 +pirs
height = 7 cm
radius = 7 cm (1/2 radius = 3.5 cm)
slant height = sqrt(((3.5)^2+(7^2))
= 7.82 cm
..
surface area of cone =( pi*(7^2)+pi*7*7.82)
=325.9 sq1.cm
TOTAl SURFACE = 325.9+307.87=633.7 sq. cm